Suppose you adjust your garden hose nozzle for a fast stream of water. You point the nozzle vertically upward at a height of 1.8 m above the ground (. When you quickly turn off the nozzle, you hear the water striking the ground next to you for another 2.6 s .


What is the water speed as it leaves the nozzle?


1
Expert's answer
2022-03-01T17:33:17-0500

"t=t_1+t_2,"

"t_1=\frac vg,"

"t_2=\frac {\sqrt{v^2+2gh}}g,"

"v=\frac{gt}2-\frac ht=12~\frac ms."